# Stochastic gravitational wave background anisotropies from inflation with non-Bunch-Davies states

### v2

@@ 3.2 Scalar-tensor-tensor bispectrum
By expanding the action Eq. (2.1) up to the cubic order in the perturbations, the interaction Hamiltonian involving one scalar and two tensors is obtained as [101]
$H_{\rm int}=-\int{\rm d}^{3}x\mathcal{L}_{shh},$ (3.8)
@@
$\qquad\qquad+b_{4}\dot{\zeta}\dot{h}_{ij}^{2}+\frac{b_{5}}{a^{2}}\partial^{2}\zeta\dot{h}_{ij}^{2}+b_{6}\psi_{,ij}\dot{h}_{ik}\dot{h}_{jk}+\frac{b_{7}}{a^{2}}\zeta_{,ij}\dot{h}_{ik}\dot{h}_{jk}\biggr{]}.$ (3.9)
As is summarized in Appendix A, the above cubic Lagrangian includes other terms proportional to the equations of motion for the linear perturbations which, in principle, can contribute to the bispectrum. However, those are always subleading under the slow-roll approximations, and hence we do not consider those in the main text.
@@
$\mathcal{B}_{shh}^{s_{2}s_{3}}$ $={\rm Re}\left[\sum^{7}_{i=1}\left(\alpha^{(s_{2})}_{k_{2}}-\beta^{(s_{2})}_{k_{2}}\right)\left(\alpha^{(s_{3})}_{k_{3}}-\beta^{(s_{3})}_{k_{3}}\right)\frac{2}{k_{1}^{3}k_{2}^{3}k_{3}^{3}}\frac{H^{6}}{\mathcal{F}_{S}\mathcal{F}_{T}^{2}c_{s}c_{h}^{2}}b_{i}\mathcal{F}_{b_{i}}\mathcal{V}_{b_{i}}+(s_{2},k_{2}\leftrightarrow s_{3},k_{3})\right].$ (3.10)
@@
$K^{\prime}$ $:=c_{s}k_{1}+c_{h}(k_{2}+k_{3}),$ (3.12)
$K_{1}^{\prime}$ $:=-c_{s}k_{1}+c_{h}(k_{2}+k_{3}),$ (3.13)
$K_{2}^{\prime}$ $:=c_{s}k_{1}+c_{h}(-k_{2}+k_{3}),$ (3.14)
$K_{3}^{\prime}$ $:=c_{s}k_{1}+c_{h}(k_{2}-k_{3}).$ (3.15)
Under Eq. (3.5) with $k_{1}=k_{L}$ and $k_{2}\simeq k_{3}\simeq k_{S}$, one can obtain
$S(\tilde{K})=\begin{cases}\displaystyle\mathcal{O}(\tilde{K}^{-(n+1)})&\ \ (\tilde{K}=K^{\prime}\ {\rm or}\ K_{1}^{\prime}),\\ \displaystyle\mathcal{O}(\eta_{0}^{(n+1)})&\ \ (\tilde{K}=K_{2}^{\prime}\ {\rm or}\ K_{3}^{\prime}).\end{cases}$ (3.16)
@@
It is worth mentioning the difference of $S(\tilde{K})$ between Eq. (3.1) and Eq. (3.5). In the former, one has $S(\tilde{K})=\mathcal{O}(\tilde{K}^{-(n+1)})$ for all of the four cases under the squeezed limit, which has resulted in the enhancements of the squeezed bispectrum in proportion to powers of $k_{S}/k_{L}$. In the latter, one has $S(\tilde{K})=\mathcal{O}(\eta_{0}^{(n+1)})$, which will result in the enhancements in proportion to powers of $-k_{S}\eta_{0}$ as will be shown below. Since we have $k_{S}/k_{L}\gg-k_{S}\eta_{0}$, the enhancements will be reduced compared to the results in the literature (e.g., Ref. [51]) where Eq. (3.1) was assumed.
@@
We summarize the details of computations in Appendix B, and we here show only the final results. The primordial STT bispectrum obtained under Eq. (3.5) is as follows,
$\mathcal{F}_{b_{1}}$ $=\frac{c_{h}^{4}k_{2}^{2}k_{3}^{2}}{H^{2}}\biggl{[}\mathcal{I}_{b_{1},K^{\prime}}\alpha^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{1},K_{1}^{\prime}}\beta^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{1},K_{2}^{\prime}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{1},K_{3}^{\prime}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}\biggr{]},$ (3.17)
$\mathcal{F}_{b_{2}}$ $=\frac{1}{H^{2}}\biggl{[}\mathcal{I}_{b_{2},K^{\prime}}\alpha^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{2},K_{1}^{\prime}}\beta^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{2},K_{2}^{\prime}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{2},K_{3}^{\prime}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}\biggr{]},$ (3.18)
$\mathcal{F}_{b_{3}}$ $=\frac{c_{h}^{2}c_{s}^{2}k_{1}^{2}k_{2}^{2}}{H^{2}}\biggl{[}\mathcal{I}_{b_{3},K^{\prime}}\alpha^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{3},K_{1}^{\prime}}\beta^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{3},K_{2}^{\prime}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{3},K_{3}^{\prime}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}\biggr{]},$ (3.19)
$\mathcal{F}_{b_{4}}$ $=\frac{c_{h}^{4}c_{s}^{2}k_{1}^{2}k_{2}^{2}k_{3}^{2}}{H}\biggl{[}\mathcal{I}_{b_{4},K^{\prime}}\alpha^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{4},K_{1}^{\prime}}\beta^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{4},K_{2}^{\prime}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{4},K_{3}^{\prime}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}\biggr{]},$ (3.20)
$\mathcal{F}_{b_{5}}$ $=c_{h}^{4}k_{1}^{2}k_{2}^{2}k_{3}^{2}\biggl{[}\mathcal{I}_{b_{5},K^{\prime}}\alpha^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{5},K_{1}^{\prime}}\beta^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{5},K_{2}^{\prime}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{5},K_{3}^{\prime}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}\biggr{]},$ (3.21)
$\mathcal{F}_{b_{6}}$ $=\mathcal{F}_{b_{4}},$ (3.22)
$\mathcal{F}_{b_{7}}$ $=\mathcal{F}_{b_{5}},$ (3.23)
@@
$\mathcal{I}_{b_{i},K^{\prime}}$ $=\mathcal{J}_{b_{i}}(k_{1},k_{2},k_{3}),$ (3.24)
$\mathcal{I}_{b_{i},K_{1}^{\prime}}$ $=-\mathcal{J}_{b_{i}}(-k_{1},k_{2},k_{3}),$ (3.25)
$\mathcal{I}_{b_{i},K_{2}^{\prime}}$ $=-\mathcal{J}_{b_{i}}(k_{1},-k_{2},k_{3})+\mathcal{S}_{b_{i}}(k_{1},-k_{2},k_{3}),$ (3.26)
$\mathcal{I}_{b_{i},K_{3}^{\prime}}$ $=-\mathcal{J}_{b_{i}}(k_{1},k_{2},-k_{3})+\mathcal{S}_{b_{i}}(k_{1},k_{2},-k_{3}),$ (3.27)
@@
$\mathcal{J}_{b_{1}}(k_{1},k_{2},k_{3})$ $:=\frac{c_{s}k_{1}+K^{\prime}}{K^{\prime 2}},$ (3.28)
$\mathcal{S}_{b_{1}}(k_{1},k_{2},k_{3})$ $:=\frac{e^{iK^{\prime}\eta_{0}}}{K^{\prime 2}}(c_{s}k_{1}+K^{\prime}-ic_{s}k_{1}K^{\prime}\eta_{0}),$ (3.29)
@@
$\qquad\qquad\quad+2c_{s}c_{h}^{2}k_{1}(k_{2}^{2}+k_{2}k_{3}+k_{3}^{2})+c_{h}^{3}(k_{2}+k_{3})(k_{2}^{2}+k_{2}k_{3}+k_{3}^{2})\biggr{]},$ (3.30)
@@
$\qquad\qquad\quad+c_{s}c_{h}k_{1}(k_{2}^{2}+4k_{2}k_{3}+k_{3}^{2}))+\frac{i}{\eta_{0}}-\frac{ic_{s}c_{h}^{2}k_{1}k_{2}k_{3}\eta_{0}}{K^{\prime}}\biggr{]},$ (3.31)
$\mathcal{J}_{b_{3}}(k_{1},k_{2},k_{3})$ $:=\frac{K^{\prime}+c_{h}k_{3}}{K^{\prime 2}},$ (3.32)
$\mathcal{S}_{b_{3}}(k_{1},k_{2},k_{3})$ $:=\frac{e^{iK^{\prime}\eta_{0}}}{K^{\prime 2}}(K^{\prime}+c_{h}k_{3}-ic_{h}K^{\prime}k_{3}\eta_{0}),$ (3.33)
$\mathcal{J}_{b_{4}}(k_{1},k_{2},k_{3})$ $:=\frac{2}{K^{\prime 3}},$ (3.34)
$\mathcal{S}_{b_{4}}(k_{1},k_{2},k_{3})$ $:=-\frac{i}{K^{\prime 3}}e^{iK^{\prime}\eta_{0}}(2i+2K^{\prime}\eta_{0}-iK^{\prime 2}\eta_{0}^{2}),$ (3.35)
$\mathcal{J}_{b_{5}}(k_{1},k_{2},k_{3})$ $:=\frac{2(3c_{s}k_{1}+K^{\prime})}{K^{\prime 4}},$ (3.36)
@@
$\qquad\qquad\quad+\frac{i\eta_{0}}{K^{\prime 3}}[K^{\prime}(-2+iK^{\prime}\eta_{0})+c_{s}k_{1}(-6+3iK^{\prime}\eta_{0}+K^{\prime 2}\eta_{0}^{2})]\biggr{\}}.$ (3.37)
For the Bunch-Davies state limit ($\beta^{(s)}_{k}\to 0$), this expression gives the previous result in Ref. [101]. In particular, in the squeezed limit the difference from the case with the Bunch-Davies state appears in $\mathcal{I}_{b_{i},K^{\prime}_{(2,3)}}$, and $\mathcal{I}_{b_{i},K^{\prime}_{(2,3)}}/\mathcal{I}_{b_{i},K^{\prime}}$ corresponds to the characteristic scale-dependent enhancement of the squeezed bispectrum with the non-Bunch-Davies states. By taking the squeezed limit ($k_{L}=k_{1}\ll k_{2}\simeq k_{3}\simeq k_{S}$), we have
$K^{\prime}\simeq K_{1}^{\prime}\simeq 2c_{h}k_{S},K_{2}^{\prime}\simeq K_{3}^{\prime}\simeq c_{s}k_{L},$ (3.38)
@@
$\mathcal{I}_{b_{1},K^{\prime}}$ $\simeq-\mathcal{I}_{b_{1},K_{1}^{\prime}}\simeq\frac{1}{2c_{h}k_{S}},\ \mathcal{I}_{b_{1},K_{2}^{\prime}}\simeq\mathcal{I}_{b_{1},K_{3}^{\prime}}\simeq i\eta_{0},$ (3.39)
$\mathcal{I}_{b_{2},K^{\prime}}$ $\simeq-\mathcal{I}_{b_{2},K_{1}^{\prime}}\simeq\frac{3c_{h}k_{S}}{2},\ \mathcal{I}_{b_{2},K_{2}^{\prime}}\simeq\mathcal{I}_{b_{2},K_{3}^{\prime}}\simeq-ic_{h}^{2}k_{S}^{2}\eta_{0},$ (3.40)
$\mathcal{I}_{b_{3},K^{\prime}}$ $\simeq-\mathcal{I}_{b_{3},K_{1}^{\prime}}\simeq\frac{3}{4c_{h}k_{S}},\ \mathcal{I}_{b_{3},K_{2}^{\prime}}\simeq-\mathcal{I}_{b_{3},K_{3}^{\prime}}\simeq\frac{c_{h}k_{S}}{2}\eta_{0}^{2},$ (3.41)
$\mathcal{I}_{b_{4},K^{\prime}}$ $\simeq-\mathcal{I}_{b_{4},K_{1}^{\prime}}\simeq\frac{1}{4c_{h}^{3}k_{S}^{3}},\ \mathcal{I}_{b_{4},K_{2}^{\prime}}\simeq\mathcal{I}_{b_{4},K_{3}^{\prime}}\simeq-\frac{i}{3}\eta_{0}^{3},$ (3.42)
$\mathcal{I}_{b_{5},K^{\prime}}$ $\simeq-\mathcal{I}_{b_{5},K_{1}^{\prime}}\simeq\frac{1}{4c_{h}^{3}k_{S}^{3}},\ \mathcal{I}_{b_{5},K_{2}^{\prime}}\simeq\mathcal{I}_{b_{5},K_{3}^{\prime}}\simeq-\frac{i}{3}\eta_{0}^{3}.$ (3.43)

### v3

@@ 3.2 STT bispectrum
By expanding the action Eq. (1) up to the cubic order in the perturbations, the interaction Hamiltonian involving one scalar and two tensors is obtained as Gao:2012ib :
$H_{\rm int}=-\int{\rm d}^{3}x\mathcal{L}_{shh},$ (47)
@@
$\qquad\qquad+b_{4}\dot{\zeta}\dot{h}_{ij}^{2}+\frac{b_{5}}{a^{2}}\partial^{2}\zeta\dot{h}_{ij}^{2}+b_{6}\psi_{,ij}\dot{h}_{ik}\dot{h}_{jk}+\frac{b_{7}}{a^{2}}\zeta_{,ij}\dot{h}_{ik}\dot{h}_{jk}\biggr].$ (48)
As is summarized in Appendix A, the above cubic Lagrangian includes other terms proportional to the equations of motion for the linear perturbations which, in principle, can contribute to the bispectrum. However, those are always subleading under the slow-roll approximations, and hence we do not consider those in the main text. Also, throughout the present paper, we approximate all coefficients of cubic interactions to constants under the slow-roll approximations.
@@
$\mathcal{B}_{shh}^{s_{2}s_{3}}$ $={\rm Re}\left[\sum^{7}_{i=1}\left(\alpha^{(s_{2})}_{k_{2}}-\beta^{(s_{2})}_{k_{2}}\right)\left(\alpha^{(s_{3})}_{k_{3}}-\beta^{(s_{3})}_{k_{3}}\right)\frac{2}{k_{1}^{3}k_{2}^{3}k_{3}^{3}}\frac{H^{6}}{\mathcal{F}_{S}\mathcal{F}_{T}^{2}c_{s}c_{h}^{2}}b_{i}\mathcal{F}_{b_{i}}\mathcal{V}_{b_{i}}+(s_{2},k_{2}\leftrightarrow s_{3},k_{3})\right].$ (49)
@@
$K^{\prime}$ $:=c_{s}k_{1}+c_{h}(k_{2}+k_{3}),$ (51)
$K_{1}^{\prime}$ $:=-c_{s}k_{1}+c_{h}(k_{2}+k_{3}),$ (52)
$K_{2}^{\prime}$ $:=c_{s}k_{1}+c_{h}(-k_{2}+k_{3}),$ (53)
$K_{3}^{\prime}$ $:=c_{s}k_{1}+c_{h}(k_{2}-k_{3}).$ (54)
Under Eq. (42) with $k_{1}=k_{L}$ and $k_{2}\simeq k_{3}\simeq k_{S}$, one can obtain
$S(\tilde{K})=\begin{cases}\displaystyle\mathcal{O}(\tilde{K}^{-(n+1)})&\ \ (\tilde{K}=K^{\prime}\ {\rm or}\ K_{1}^{\prime}),\\ \displaystyle\mathcal{O}(\eta_{0}^{(n+1)})&\ \ (\tilde{K}=K_{2}^{\prime}\ {\rm or}\ K_{3}^{\prime}).\end{cases}$ (55)
@@
It is worth mentioning the difference of S(K~)S($\tilde${K}) between Eq. (36) and Eq. (42). In the former, one has $S(\tilde{K})=\mathcal{O}(\tilde{K}^{-(n+1)})$ for all four cases under the squeezed limit, which has resulted in the enhancements of the squeezed bispectrum being in proportion to powers of $k_{S}/k_{L}$. In the latter, one has $S(\tilde{K})=\mathcal{O}(\eta_{0}^{(n+1)})$, which will result in the enhancements being in proportion to powers of $k_{S}\eta_{0}$ as will be shown below. Since we have $k_{S}/k_{L}\gg-k_{S}\eta_{0}$, the enhancements will be reduced compared to the results in the literature (e.g. Ref. Ganc:2012ae ) where Eq. (36) was assumed.
@@
We summarize the computational details in Appendix B, and we here show only the final results. The primordial STT bispectrum obtained under Eq. (42) is as follows,
$\mathcal{F}_{b_{1}}$ $=\frac{c_{h}^{4}k_{2}^{2}k_{3}^{2}}{H^{2}}\biggl[\mathcal{I}_{b_{1},K^{\prime}}\alpha^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{1},K_{1}^{\prime}}\beta^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{1},K_{2}^{\prime}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{1},K_{3}^{\prime}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}\biggr$], (56)
$\mathcal{F}_{b_{2}}$ $=\frac{1}{H^{2}}\biggl[\mathcal{I}_{b_{2},K^{\prime}}\alpha^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{2},K_{1}^{\prime}}\beta^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{2},K_{2}^{\prime}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{2},K_{3}^{\prime}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}\biggr$], (57)
$\mathcal{F}_{b_{3}}$ $=\frac{c_{h}^{2}c_{s}^{2}k_{1}^{2}k_{2}^{2}}{H^{2}}\biggl[\mathcal{I}_{b_{3},K^{\prime}}\alpha^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{3},K_{1}^{\prime}}\beta^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{3},K_{2}^{\prime}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{3},K_{3}^{\prime}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}\biggr$], (58)
$\mathcal{F}_{b_{4}}$ $=\frac{c_{h}^{4}c_{s}^{2}k_{1}^{2}k_{2}^{2}k_{3}^{2}}{H}\biggl[\mathcal{I}_{b_{4},K^{\prime}}\alpha^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{4},K_{1}^{\prime}}\beta^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{4},K_{2}^{\prime}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{4},K_{3}^{\prime}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}\biggr$], (59)
$\mathcal{F}_{b_{5}}$ $=c_{h}^{4}k_{1}^{2}k_{2}^{2}k_{3}^{2}\biggl[\mathcal{I}_{b_{5},K^{\prime}}\alpha^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}^{*}_{b_{5},K_{1}^{\prime}}\beta^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{5},K_{2}^{\prime}}\beta^{(s_{2})*}_{k_{2}}\alpha^{(s_{3})*}_{k_{3}}+\mathcal{I}_{b_{5},K_{3}^{\prime}}\alpha^{(s_{2})*}_{k_{2}}\beta^{(s_{3})*}_{k_{3}}\biggr$], (60)
$\mathcal{F}_{b_{6}}$ $=\mathcal{F}_{b_{4}},$ (61)
$\mathcal{F}_{b_{7}}$ $=\mathcal{F}_{b_{5}},$ (62)
@@
$\mathcal{I}_{b_{i},K^{\prime}}$ $=\mathcal{J}_{b_{i}}(k_{1},k_{2},k_{3}),$ (63)
$\mathcal{I}_{b_{i},K_{1}^{\prime}}$ $=-\mathcal{J}_{b_{i}}(-k_{1},k_{2},k_{3}),$ (64)
$\mathcal{I}_{b_{i},K_{2}^{\prime}}$ $=-\mathcal{J}_{b_{i}}(k_{1},-k_{2},k_{3})+\mathcal{S}_{b_{i}}(k_{1},-k_{2},k_{3}),$ (65)
$\mathcal{I}_{b_{i},K_{3}^{\prime}}$ $=-\mathcal{J}_{b_{i}}(k_{1},k_{2},-k_{3})+\mathcal{S}_{b_{i}}(k_{1},k_{2},-k_{3}),$ (66)
@@
$\mathcal{J}_{b_{1}}(k_{1},k_{2},k_{3})$ $:=\frac{c_{s}k_{1}+K^{\prime}}{K^{\prime 2}},$ (67)
$\mathcal{S}_{b_{1}}(k_{1},k_{2},k_{3})$ $:=\frac{e^{iK^{\prime}\eta_{0}}}{K^{\prime 2}}(c_{s}k_{1}+K^{\prime}-ic_{s}k_{1}K^{\prime}\eta_{0}),$ (68)
@@
$\qquad\qquad\quad+2c_{s}c_{h}^{2}k_{1}(k_{2}^{2}+k_{2}k_{3}+k_{3}^{2})+c_{h}^{3}(k_{2}+k_{3})(k_{2}^{2}+k_{2}k_{3}+k_{3}^{2})\biggr],$ (69)
@@
$\qquad\qquad\quad+c_{s}c_{h}k_{1}(k_{2}^{2}+4k_{2}k_{3}+k_{3}^{2}))+\frac{i}{\eta_{0}}-\frac{ic_{s}c_{h}^{2}k_{1}k_{2}k_{3}\eta_{0}}{K^{\prime}}\biggr],$ (70)
$\mathcal{J}_{b_{3}}(k_{1},k_{2},k_{3})$ $:=\frac{K^{\prime}+c_{h}k_{3}}{K^{\prime 2}},$ (71)
$\mathcal{S}_{b_{3}}(k_{1},k_{2},k_{3})$ $:=\frac{e^{iK^{\prime}\eta_{0}}}{K^{\prime 2}}(K^{\prime}+c_{h}k_{3}-ic_{h}K^{\prime}k_{3}\eta_{0}),$ (72)
$\mathcal{J}_{b_{4}}(k_{1},k_{2},k_{3})$ $:=\frac{2}{K^{\prime 3}},$ (73)
$\mathcal{S}_{b_{4}}(k_{1},k_{2},k_{3})$ $:=-\frac{i}{K^{\prime 3}}e^{iK^{\prime}\eta_{0}}(2i+2K^{\prime}\eta_{0}-iK^{\prime 2}\eta_{0}^{2}),$ (74)
$\mathcal{J}_{b_{5}}(k_{1},k_{2},k_{3})$ $:=\frac{2(3c_{s}k_{1}+K^{\prime})}{K^{\prime 4}},$ (75)
@@
$\qquad\qquad\quad+\frac{i\eta_{0}}{K^{\prime 3}}[K^{\prime}(-2+iK^{\prime}\eta_{0})+c_{s}k_{1}(-6+3iK^{\prime}\eta_{0}+K^{\prime 2}\eta_{0}^{2})]\biggr\}.$ (76)
For the Bunch-Davies state limit ($\beta^{(s)}_{k}\to 0$), this expression gives the previous result in Ref. Gao:2012ib . In particular, in the squeezed limit the difference from the case with the Bunch-Davies state appears in $\mathcal{I}_{b_{i},K^{\prime}_{(2,3)}}$, and $\mathcal{I}_{b_{i},K^{\prime}_{(2,3)}}/\mathcal{I}_{b_{i},K^{\prime}}$ corresponds to the characteristic scale-dependent enhancement of the squeezed bispectrum with the non-Bunch-Davies states. By taking the squeezed limit ($k_{L}=k_{1}\ll k_{2}\simeq k_{3}\simeq k_{S}$), we have
$K^{\prime}\simeq K_{1}^{\prime}\simeq 2c_{h}k_{S},K_{2}^{\prime}\simeq K_{3}^{\prime}\simeq c_{s}k_{L},$ (77)
@@
$\mathcal{I}_{b_{1},K^{\prime}}$ $\simeq-\mathcal{I}_{b_{1},K_{1}^{\prime}}\simeq\frac{1}{2c_{h}k_{S}},\ \mathcal{I}_{b_{1},K_{2}^{\prime}}\simeq\mathcal{I}_{b_{1},K_{3}^{\prime}}\simeq i\eta_{0},$ (78)
$\mathcal{I}_{b_{2},K^{\prime}}$ $\simeq-\mathcal{I}_{b_{2},K_{1}^{\prime}}\simeq\frac{3c_{h}k_{S}}{2},\ \mathcal{I}_{b_{2},K_{2}^{\prime}}\simeq\mathcal{I}_{b_{2},K_{3}^{\prime}}\simeq-ic_{h}^{2}k_{S}^{2}\eta_{0},$ (79)
$\mathcal{I}_{b_{3},K^{\prime}}$ $\simeq-\mathcal{I}_{b_{3},K_{1}^{\prime}}\simeq\frac{3}{4c_{h}k_{S}},\ \mathcal{I}_{b_{3},K_{2}^{\prime}}\simeq-\mathcal{I}_{b_{3},K_{3}^{\prime}}\simeq\frac{c_{h}k_{S}}{2}\eta_{0}^{2},$ (80)
$\mathcal{I}_{b_{4},K^{\prime}}$ $\simeq-\mathcal{I}_{b_{4},K_{1}^{\prime}}\simeq\frac{1}{4c_{h}^{3}k_{S}^{3}},\ \mathcal{I}_{b_{4},K_{2}^{\prime}}\simeq\mathcal{I}_{b_{4},K_{3}^{\prime}}\simeq-\frac{i}{3}\eta_{0}^{3},$ (81)
$\mathcal{I}_{b_{5},K^{\prime}}$ $\simeq-\mathcal{I}_{b_{5},K_{1}^{\prime}}\simeq\frac{1}{4c_{h}^{3}k_{S}^{3}},\ \mathcal{I}_{b_{5},K_{2}^{\prime}}\simeq\mathcal{I}_{b_{5},K_{3}^{\prime}}\simeq-\frac{i}{3}\eta_{0}^{3}.$ (82)